Circular No. 8938

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 H1 (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(discovery observation below) and posted on the Minor Planet Center’s ‘NEOCP’ webpage, was first reported as cometary by E. Reina L., Hospitalet, Spain, on Apr. 19.0 UT (0.25-m f/3.3 Schmidt-Cassegrain reflector), who noted a 25′′ coma of total mag 16.6 and a 43′′ tail in p.a. 233^{o}. E. Guido and G. Sostero, observing remotely from Mayhill, NM, U.S.A., on Apr. 19.5 (0.25-m f/3.4 reflector), report that stacking twenty unfiltered 60-s CCD exposures in strong moonlight showed the object to be diffuse in comparison to stars of similar brightness. J. M. Aymami, Tiana, Spain (0.13-m f/5.9 refractor), remarked on a suspected coma in p.a. 209^{o} on Apr. 19.8. At the same time R. Apitzsch, Wildberg, Germany (0.35-m f/4.2 reflector) noted a tail to the southwest.

2008 UT  α_{2000}  δ_{2000}  Mag.
Apr. 18.34722  17^{h}02^{m}01.^{s}68  +79^{o}15′56′′.0  18.1

The available astrometry, the preliminary parabolic orbital elements (T = 2008 Mar. 11.781 TT, q = 2.78957 AU, ω = 94^{o}.644, Ω = 33^{o}.724, i = 75^{o}.770, equinox 2000.0), and an ephemeris appear on MPEC 2008-H09.

V2491 CYGNI R. J. Rudy, D. K. Lynch, and R. W. Russell, The Aerospace Corpora-

tion; C. E. Woodward, University of Minnesota; and K. Covey, Center for Astrophysics, report on SpeX observations (wavelength range 0.8–2.5 μm) of this nova (cf. IAUC 8934) obtained at the Infrared Telescope Facility on Apr. 17.6 UT. In the five days since Lynch et al. (IAUC 8935) observed V2491 Cyg on Apr. 12, it declined in brightness by a factor of 3. The strengths of the emission lines compared to the continuum have generally increased, as has the excitation of the emission-line spectrum. The neutral helium triplet and singlet at 1.08 and 2.06 μm, respectively, doubled in size relative to the hydrogen lines. Although the C I and N I lines decreased, the O I lines at 0.84 and 1.13 μm, which are fluorescently excited by Lyβ, are now the strongest emission features in the infrared spectrum. Their increase has permitted more accurate measurements of their relative strengths, which indicate a reddening of E(B–V) = 0.43, a value slightly greater than that reported by Lynch et al. Photometric magnitudes on Apr. 17: J = 7.7, H = 7.8, K = 7.4.

2008 April 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green